Find the equation of the bisector of the acute angles and also the equation of the bisector of
the obtuse angles formed by the lines
7x − 24y = 8
and
3x + 4y = 12.

Answer by Alan3354(69443)

Find the equation of the bisector of the acute angles and also the equation of the bisector of
the obtuse angles formed by the lines
7x − 24y = 8
and
3x + 4y = 12.
-----------------------
Find the slope of each line, and the intersection:
L1: 7x − 24y = 8 - m1 = 7/24
L2: 3x + 4y = 12 - m2 = -3/4
==========
7x − 24y = 8
18x + 24y = 72
---------------------------- Add
25x = 80
x = 16/5
y = 3/5
Intersection (16/5,3/5)
==========================================
The slope is the tangent of the angle with the x-axis.
L1: Angle1 = 16.26 degs
l2: Angle2 = -36.87 degs
(16.26 - 36.87)/2 = -10.3 degs
Slope of bisector = tan(-10.3) = -0.181818... = -2/11
y-0.6 = (-2/11)*(x - 3.2) is the bisector of the acute angle
=============
The slope of the other bisector is the negative inverse, = 11/2
y-0.6 = (11/2)*(x - 3.2) is the bisector of the obtuse angle